The floor of a square room is covered in square tiles. If the diagonals of the room have a total of 21 tiles, how many tiles are on the floor?
Answer by nerdybill(5404)

The floor of a square room is covered in square tiles. If the diagonals of the room have a total of 21 tiles, how many tiles are on the floor?
.
Let x = number of tiles on each side of the square room
then applying Pythagorean theorem:
x^2 + x^2 = 21^2
2x^2 = 441
x^2 = 220.5
x = 14.85 tiles on each side
.
Total number of tiles:
x*x = x^2 = 14.85^2 = 220.5 tiles